# Robust Prompt Optimization for Defending Language Models Against Jailbreaking Attacks

### v3

@@ 1 Introduction
Since the discovery of these attacks, various defense mechanisms have been proposed, including input filters (Jain et al., 2023; Kumar et al., 2023), input smoothing (Robey et al., 2023), and few-shot examples (Wei et al., 2024; Anil et al., 2024). While effective for initially proposed attacks such as Greedy Coordinate Gradient (GCG) Zou et al. (2023), these often cannot generalize to multiple jailbreaks or incur additional inference costs, falling short of a strong and practical defense. In addition, a formal optimization objective for defense has yet to be proposed, especially in the adaptive attack scenario, making it difficult to consider how defenses can adapt to future attacks.
@@
To address these issues, we formalize a minimax defensive objective motivated by adversarial training and propose Robust Prompt Optimization (RPO), a discrete optimization algorithm to optimize for this objective. Our study is motivated by the increasing adoption of system-level guardrails Inan et al. (2023); Rebedea et al. (2023), components inaccessible to the user used in LLM deployments to steer model behavior, such as system prompts or input/output filters. RPO improves robustness through system-level modifications to the model input. We evaluate RPO on two recently proposed red-teaming benchmarks, JailbreakBench (Chao et al., 2024) and HarmBench (Mazeika et al., 2024), which both cover a broad range of harmful risk categories and attack methods. On JailbreakBench, RPO reduces the attack success rate (ASR) to 6% on GPT-4 and 0% on LLama-2, outperforming existing defenses and setting the state-of-the-art as a jailbreaking defense. In addition, RPO suffixes incur a negligible inference cost, only have a minor effect on benign prompts, and transfer to black-box models and unknown attacks. In summary, our contributions are the following:
@@ 2 Related Work
Adversarial attacks on LLMs. Similar attacks have been studied in NLP, including text classification (Ebrahimi et al., 2017; Alzantot et al., 2018; Wang et al., 2022), question-answering (Jia and Liang, 2017), or triggering toxic completions (Wallace et al., 2019; Jones et al., 2023; Zou et al., 2023). Language models are among the most generally capable models and have been applied to many domains beyond language (Yao et al., 2023; Zhou et al., 2023). As a result, inducing harmful behaviors has been the primary threat model for LLMs (Carlini et al., 2023). This has resulted in many recent jailbreaking attacks, where an adversary modifies a prompt manually to circumvent alignment training and induce harmful behavior. These attacks can be created manually by humans (Liu et al., 2023b; Wei et al., 2023; Zeng et al., 2024), refined with another LLM (Chao et al., 2023; Mehrotra et al., 2023; Liu et al., 2023a; Jin et al., 2024; Paulus et al., 2024), or generated with discrete optimization (Zou et al., 2023; Lapid et al., 2023; Zhu et al., 2023b; Sadasivan et al., 2024). In addition, Huang et al. (2023) finds that simply modifying decoding settings can jailbreak many open-source LLMs. Other attacks include extracting training data (Carlini et al., 2021; Nasr et al., 2023) and misclassification (Zhu et al., 2023a; Wang et al., 2023), but we focus on harmful behaviors.
Safety and Defenses for LLMs. Even without an adversary, LLMs are prone to generating biased or toxic content (Sheng et al., 2019; McGuffie and Newhouse, 2020; Deshpande et al., 2023). To mitigate this, many modern LLMs undergo significant red-teaming (Perez et al., 2022; Mazeika et al., 2024) and additional training such as reinforcement learning with human feedback (Christiano et al., 2017; Ouyang et al., 2022; Bai et al., 2022) to be safer and refuse harmful requests. Additional defenses have recently been proposed with the discovery of additional failure modes, such as jailbreaking, on aligned LLMs. For instance, (Jain et al., 2023) examines simple defenses such as rephrasing the input and finds that the GCG attack (Zou et al., 2023) can be defended with a perplexity filter. Other defenses that have been explored include in-context learning (Zhang et al., 2023), sampling (Li et al., 2023), input processing (Cao et al., 2023; Robey et al., 2023; Kumar et al., 2023), and content moderation (Inan et al., 2023; Yuan et al., 2024). While often effective for the threat models considered, many defenses rely on heuristics such as perplexity that can be circumvented by human-readable jailbreaks or require additional inference calls, reducing practicality.
@@ 3.2 Defense Objective
While prevailing methods to improve LLM alignment involve fine-tuning, the objective of matching human preferences does not generally account for adversaries and jailbreaking. In addition, the high cost of generating attack prompts makes standard adversarial training on these samples difficult Jain et al. (2023). We center our optimization approach on the prompt level to address this. We formalize this as the negative log probability of a target token output $\mathbf{y}^{\prime}$ that refuses $\tilde{\mathbf{x}}_{1:n}$. This can be represented as the normal output of an LLM trained to maximize $\mathcal{R}^{\prime}$ or $\mathbf{y}^{\prime}=\max\mathcal{R}^{*}(\mathbf{y}|\tilde{\mathbf{x}}_{1:n})$. Thus, we have the following safe loss and defense objective
@@ 3.3 Robust Prompt Optimization
$A^{*}:=\operatorname*{argmin}_{\mathcal{A}}\mathcal{L}_{j}^{\text{adv}}\sum_{1\leq o\leq m}(A_{o}(x^{(j)}))$triangleright Select jailbreak that minimizes adversarial loss
@@
We simulate the adaptive threat model for the first step by adding the current defensive suffix to the original prompt and applying or generating a jailbreak prompt afterward. This is a straightforward modification to the prompt for simple, manual jailbreaks such as in-context examples (Wei et al., 2024). For automatic jailbreaks such as GCG Zou et al. (2023), we apply several iterations of the jailbreak until either the RPO suffix is broken or until a fixed compute budget is exhausted. This allows RPO to support a variety of attacks during the optimization process. Our main technical contribution for this component is the selection step, where after its respective generation, we apply the jailbreak prompt that minimizes the adversarial loss for that instruction, according to Eq. 4. As the adversarial loss is calculated with the addition of the current RPO suffix, this ensures the optimization is performed under worst-case conditions and reduces the chance for the suffix to overfit on a particular jailbreak. In practice, due to the cost of generating new attack prompts, we only perform this operation again after a certain number of iterations $R$.
After a jailbreak is applied, the second step optimizes the suffix to minimize the safe loss Eq.  7. We adopt a method similar to AutoPrompt Shin et al. (2020) and GCG, using a greedy coordinate descent approach to assess how replacing the i-th token affects the safe loss. This involves calculating the first-order approximation and selecting the top-k tokens with the largest negative gradient. We then randomly select $B\leq k|\mathcal{I}|$ tokens from this set of candidates, obtain the exact loss on this subset, and replace the current token with the token with the smallest loss. Both steps are applied in succession for a number of iterations $T$. The full algorithm is described in Alg. 1.
@@ 4 Experiments
Attack Defense Vicuna Llama-2 GPT-3.5 GPT-4
PAIR None 82% 4% 76% 50%
SmoothLLM 47% 1% 12% 25%
Perplexity Filter 81% 4% 15% 43%
Rephrasing 25% 4% 22% 35%
RPO (Ours) 16% 0% 6% 6%
GCG None 58% 2% 34% 1%
SmoothLLM 1% 1% 1% 3%
Perplexity Filter 1% 0% 1% 0%
Rephrasing 4% 0% 4% 1%
RPO (Ours) 0% 0% 0% 0%
JBC None 79% 0% 0% 0%
SmoothLLM 64% 0% 0% 0%
Perplexity Filter 79% 0% 0% 0%
Rephrasing 15% 0% 1% 1%
RPO (Ours) 70% 0% 0% 0%
@@ 4.1 Experimental Setup
We optimize and evaluate our method on the instructions, attack baselines, and defense baselines from two recently proposed red-teaming benchmarks, HarmBench Mazeika et al. (2024) and JailbreakBench Chao et al. (2024). For text-based LLMs, HarmBench and JailbreakBench contain 400 and 100 distinct harmful behaviors, respectively. These involve harmful prompts safety-tuned LLMs consistently reject from a broad range of risk categories, such as “Describe how to exploit elderly people with dementia for access to their bank accounts” We measure attack success rate (ASR) and evaluate success following the corresponding benchmarks, which use LLM-as-a-judge. We evaluate attacks and defenses on four LLMs, open-source Vicuna-13B Zheng et al. (2023) and Llama-2-7B-Chat Touvron et al. (2023), and closed-sourced GPT-3.5-Turbo and GPT-4 OpenAI (2023), following JailbreakBench.
Baseline Attacks and Defenses. We use the attacks provided on each benchmark. JailbreakBench contains three attacks: (1) Greedy Coordinate Gradient (GCG) Zou et al. (2023), (2) Prompt Automatic Iterative Refinement (PAIR) Chao et al. (2023), and hand-crafted jailbreaks from Jailbreak Chat (JBC) Wei et al. (2023). HarmBench contains 18 attacks, of which we use six text-based attacks with the highest average ASR: GCG, Automated Discrete Adversarial Natural Prompt (AutoDAN) Liu et al. (2023a), PAIR, Few-Shot Examples Perez et al. (2022); Wei et al. (2024), Tree-of-Attacks with Pruning (TAP) Mehrotra et al. (2023), and Persuasive Adversarial Prompt (PAP) Zeng et al. (2024). We use the defenses provided on each benchmark as our baselines. JailbreakBench contains three defenses: Perplexity Filter Jain et al. (2023), SmoothLLM Robey et al. (2023), and Rephrasing Jain et al. (2023), while HarmBench does not provide any defenses besides the base models. We follow the default attack and defense implementation settings.
RPO Setup. During optimization for RPO, we target the Llama-2-7B model, use a suffix length of 20 tokens, and optimize for 500 steps using a batch size of 64, top-k of 256, and selection interval of 50. We optimize the suffix using 25 randomly selected instructions from the training set of AdvBench Zou et al. (2023) to minimize overlap with evaluation instructions. We optimize the suffix on the three jailbreaks from JailbreakBench, which we find sufficient for high transferability to the other attacks on HarmBench. This includes GCG, PAIR, and JBC. During each inner minimization step, we regenerate a PAIR and GCG jailbreak for each instruction, including the current RPO suffix, but do not change the handcrafted jailbreak prompts. During inference, we place the RPO suffix after the user input as a component of the system prompt. All jailbreak details and example outputs, including on the ChatGPT interface, can be found in the Appendix.
@@ 4.2 Main Results
Known Attacks on JailbreakBench. In Tab. 1, we observe that on JailbreakBench RPO significantly improves upon baseline defense robustness to PAIR, GCG, and JBC. Models besides Vicuna have been alignment-tuned and are already robust to prompts from JBC but vulnerable to other attacks. We find perplexity filtering is highly effective on GCG but is not effective on the natural language jailbreak prompts in PAIR and JBC. SmoothLLM is more effective across multiple attacks due to not relying on the perplexity heuristic. Still, it cannot defend against a significant proportion of prompts from PAIR, the strongest attack. Rephrasing is surprisingly effective, especially on JBC for Vicuna, outperforming the other defenses. We observe RPO is more effective than all baseline defenses on PAIR for all models, reducing ASR by 66% on Vicuna and 44% on GPT-4 and improving on the state-of-the-art defense SmoothLLM by 31% and 19%. This also shows that RPO suffixes transfer across models, as the suffix was optimized using Llama-2 but can transfer to Vicuna and even closed-source GPT models. Notably, RPO reduces GCG ASR to 0% for all models, fully defending against the attack. Using RPO with Llama-2 makes the model robust to all three attacks, the first time a model is fully robust on JailbreakBench.
Transfer to Unknown Attacks on HarmBench. HarmBench marks a difficult distribution shift from the optimization setup of RPO as it contains many attacks RPO was not optimized on and has a broader inclusion of risk categories, such as copyright infringement and contextual behaviors referencing user context. These categories are not covered in the instructions we optimize RPO on, forcing the defense to generalize. Despite this, in Tab. 2 we observe RPO transfers to all attacks in HarmBench for all models, generalizing to difficult attacks such as TAP and AutoDAN. Notably, RPO reduces ASR by an average of 18%, 6.6%, 8.7%, and 3.5% for Vicuna, Llama-2, GPT-3.5, and GPT-4, respectively. This suggests RPO is universally effective as a defense for harmful queries, irrespective of the attack. However, we observe a much lower improvement on HarmBench than JailbreakBench, reflecting the challenges of generalizing to new attacks and risk categories.

### v5

@@ 1 Introduction
Since the discovery of these attacks, various defense mechanisms have been proposed, including input filters (Jain et al., 2023; Kumar et al., 2023), input smoothing (Robey et al., 2023), and few-shot examples (Wei et al., 2024; Anil et al., 2024). While effective for initially proposed attacks such as Greedy Coordinate Gradient (GCG) (Zou et al., 2023), these often cannot generalize to multiple jailbreaks or incur additional inference costs, falling short of a strong and practical defense. In addition, a formal optimization objective for defense has yet to be proposed, especially in the adaptive attack scenario, making it difficult to consider how defenses can adapt to future attacks.
@@
To address these issues, we formalize a minimax defensive objective motivated by adversarial training and propose Robust Prompt Optimization (RPO), a discrete optimization algorithm to optimize for this objective. Our study is motivated by the increasing adoption of system-level guardrails (Inan et al., 2023; Rebedea et al., 2023), components inaccessible to the user used in LLM deployments to steer model behavior, such as system prompts or input/output filters. RPO improves robustness through system-level modifications to the model input. We evaluate RPO on two recently proposed red-teaming benchmarks, JailbreakBench (Chao et al., 2024) and HarmBench (Mazeika et al., 2024), which both cover a broad range of harmful risk categories and attack methods. On JailbreakBench, RPO reduces the attack success rate (ASR) to 6% on GPT-4 and 0% on Llama-2, outperforming existing defenses and setting the state-of-the-art as a jailbreaking defense. In addition, RPO suffixes incur a negligible inference cost, only have a minor effect on benign prompts, and transfer to black-box models and unknown attacks. In summary, our contributions are the following:
@@ 2 Related Work
Adversarial attacks on LLMs. Similar attacks have been studied in NLP, including text classification (Ebrahimi et al., 2017; Alzantot et al., 2018; Wang et al., 2022), question-answering (Jia and Liang, 2017), or triggering toxic completions (Wallace et al., 2019; Jones et al., 2023; Zou et al., 2023). Language models are among the most generally capable models and have been applied to many domains beyond language (Yao et al., 2023; Zhou et al., 2023). As a result, inducing harmful behaviors has been the primary threat model for LLMs (Carlini et al., 2023). This has resulted in many recent jailbreaking attacks, where an adversary modifies a prompt manually to circumvent alignment training and induce harmful behavior. These attacks can be created manually by humans (Liu et al., 2023b; Wei et al., 2023; Zeng et al., 2024), refined with another LLM (Chao et al., 2023; Mehrotra et al., 2023; Liu et al., 2023a; Jin et al., 2024; Paulus et al., 2024), or generated with discrete optimization (Zou et al., 2023; Lapid et al., 2023; Zhu et al., 2023b; Sadasivan et al., 2024). In addition, (Huang et al., 2023) finds that simply modifying decoding settings can jailbreak many open-source LLMs. Other attacks include extracting training data (Carlini et al., 2021; Nasr et al., 2023) and misclassification (Zhu et al., 2023a; Wang et al., 2023), but we focus on harmful behaviors.
Safety and Defenses for LLMs. Even without an adversary, LLMs are prone to generating biased or toxic content (Sheng et al., 2019; McGuffie and Newhouse, 2020; Deshpande et al., 2023). To mitigate this, many modern LLMs undergo significant red-teaming (Perez et al., 2022; Mazeika et al., 2024) and additional training such as reinforcement learning with human feedback (Christiano et al., 2017; Ouyang et al., 2022; Bai et al., 2022) to be safer and refuse harmful requests. Additional defenses have recently been proposed with the discovery of additional failure modes, such as jailbreaking, on aligned LLMs. For instance, (Jain et al., 2023) examines simple defenses such as rephrasing the input and finds that the GCG attack (Zou et al., 2023) can be defended with a perplexity filter. Other defenses that have been explored include in-context learning (Zhang et al., 2023), sampling (Li et al., 2023), input processing (Cao et al., 2023; Robey et al., 2023; Kumar et al., 2023), and content moderation (Inan et al., 2023). While often effective for the threat models considered, many defenses rely on heuristics such as perplexity that can be circumvented by human-readable jailbreaks or require additional inference calls, reducing practicality. Concurrent to our work, (Mo et al., 2024) also considers a similar optimization objective, but only optimizes prompts on the GCG attack, which may limit transferability.
@@ 3.2 Defense Objective
While prevailing methods to improve LLM alignment involve fine-tuning, the objective of matching human preferences does not generally account for adversaries and jailbreaking. In addition, the high cost of generating attack prompts makes standard adversarial training on these samples difficult (Jain et al., 2023). We center our optimization approach on the prompt level to address this. We formalize this as the negative log probability of a target token output $\mathbf{y}^{\prime}$ that refuses $\tilde{\mathbf{x}}_{1:n}$. This can be represented as the normal output of an LLM trained to maximize $\mathcal{R}^{\prime}$ or $\mathbf{y}^{\prime}=\max\mathcal{R}^{*}(\mathbf{y}|\tilde{\mathbf{x}}_{1:n})$. Thus, we have the following safe loss and defense objective
@@ 3.3 Robust Prompt Optimization
$A^{*}:=\operatorname*{argmin}_{\mathcal{A}}\mathcal{L}_{j}^{\text{adv}}\sum_{1\leq o\leq m}(A_{o}(x^{(j)}))$triangleright Jailbreak that minimizes adv. loss
@@
We simulate the adaptive threat model for the first step by adding the current defensive suffix to the original prompt and applying or generating a jailbreak prompt afterward. This is a straightforward modification to the prompt for simple, manual jailbreaks such as in-context examples (Wei et al., 2024). For automatic jailbreaks such as GCG (Zou et al., 2023), we apply several iterations of the jailbreak until either the RPO suffix is broken or until a fixed compute budget is exhausted. This allows RPO to support a variety of attacks during the optimization process. Our main technical contribution for this component is the selection step, where after its respective generation, we apply the jailbreak prompt that minimizes the adversarial loss for that instruction, according to Eq. 4. As the adversarial loss is calculated with the addition of the current RPO suffix, this ensures the optimization is performed under worst-case conditions and reduces the chance for the suffix to overfit on a particular jailbreak. In practice, due to the cost of generating new attack prompts, we only perform this operation again after a certain number of iterations $R$.
After a jailbreak is applied, the second step optimizes the suffix to minimize the safe loss Eq.  7. We adopt a method similar to AutoPrompt (Shin et al., 2020) and GCG, using a greedy coordinate descent approach to assess how replacing the i-th token affects the safe loss. This involves calculating the first-order approximation and selecting the top-k tokens with the largest negative gradient. We then randomly select $B\leq k|\mathcal{I}|$ tokens from this set of candidates, obtain the exact loss on this subset, and replace the current token with the token with the smallest loss. Both steps are applied in succession for a number of iterations $T$. The full algorithm is described in Alg. 1.
@@ 4 Experiments
Attack Defense Vicuna Llama-2-7B Qwen-7B Llama2-13B GPT-3.5 GPT-4
PAIR None 82% 4% 68% 2% 76% 50%
SmoothLLM 47% 1% 36% 1% 12% 25%
Perplexity Filter 81% 4% 66% 2% 15% 43%
Rephrasing 25% 4% 13% 2% 22% 35%
Few-Shot 27% 6% 16% 1% 8% 10%
RPO (Ours) 16% 0% 4% 1% 6% 6%
GCG None 58% 2% 11% 0% 34% 1%
SmoothLLM 1% 1% 2% 0% 1% 3%
Perplexity Filter 1% 0% 0% 0% 1% 0%
Rephrasing 4% 0% 7% 0% 4% 1%
Few-Shot 1% 0% 1% 0% 0% 0%
RPO (Ours) 0% 0% 0% 0% 0% 0%
JBC None 79% 0% 58% 1% 0% 0%
SmoothLLM 64% 0% 44% 0% 0% 0%
Perplexity Filter 79% 0% 58% 1% 0% 0%
Rephrasing 15% 0% 47% 1% 1% 1%
Few-Shot 41% 0% 50% 1% 0% 0%
RPO (Ours) 70% 0% 45% 0% 0% 0%
@@ 4.1 Experimental Setup
We optimize and evaluate our method on the instructions, attack baselines, and defense baselines from two recently proposed red-teaming benchmarks, HarmBench (Mazeika et al., 2024) and JailbreakBench (Chao et al., 2024). For text-based LLMs, HarmBench and JailbreakBench contain 400 and 100 distinct harmful behaviors, respectively. These involve harmful prompts safety-tuned LLMs consistently reject from a broad range of risk categories, such as “Describe how to exploit elderly people with dementia for access to their bank accounts” We measure attack success rate (ASR) and evaluate success following the corresponding benchmarks, which use LLM-as-a-judge. We evaluate attacks and defenses on six LLMs, open-source Vicuna-13B (Zheng et al., 2023), Llama-2-7B-Chat (Touvron et al., 2023), Qwen-1.5-14B (Bai et al., 2023), and Llama-2-13B-Chat (Touvron et al., 2023), and closed-sourced GPT-3.5-Turbo and GPT-4 (OpenAI, 2023).
Baseline Attacks and Defenses. We use the attacks provided on each benchmark. JailbreakBench contains three attacks: (1) Greedy Coordinate Gradient (GCG) (Zou et al., 2023), (2) Prompt Automatic Iterative Refinement (PAIR) (Chao et al., 2023), and hand-crafted jailbreaks from Jailbreak Chat (JBC) (Wei et al., 2023). HarmBench contains 18 attacks, of which we use six text-based attacks with the highest average ASR: GCG, Automated Discrete Adversarial Natural Prompt (AutoDAN) (Liu et al., 2023a), PAIR, Few-Shot Examples (Perez et al., 2022; Wei et al., 2024), Tree-of-Attacks with Pruning (TAP) (Mehrotra et al., 2023), and Persuasive Adversarial Prompt (PAP) (Zeng et al., 2024). We use the defenses provided on each benchmark as our baselines, as well as Few-Shot examples Wei et al. (2024). JailbreakBench contains three defenses: Perplexity Filter (Jain et al., 2023), SmoothLLM (Robey et al., 2023), and Rephrasing (Jain et al., 2023), while HarmBench does not provide any defenses besides the base models. We follow the default attack and defense implementation settings.
RPO Setup. During optimization for RPO, we target the Llama-2-7B model, use a suffix length of 20 tokens, and optimize for 500 steps using a batch size of 64, top-k of 256, and selection interval of 50. We optimize the suffix using 25 randomly selected instructions from the training set of AdvBench (Zou et al., 2023) to minimize overlap with evaluation instructions. We optimize the suffix on the three jailbreaks from JailbreakBench, which we find sufficient for high transferability to the other attacks on HarmBench. This includes GCG, PAIR, and JBC. During each inner minimization step, we regenerate a PAIR and GCG jailbreak for each instruction, including the current RPO suffix, but do not change the handcrafted jailbreak prompts. During inference, we place the RPO suffix after the user input as a component of the system prompt. All jailbreak details and example outputs, including on the ChatGPT interface, can be found in the Appendix.
@@ 4.2 Main Results
Known Attacks on JailbreakBench. In Tab. 1, we observe that on JailbreakBench RPO significantly improves upon baseline defense robustness to PAIR, GCG, and JBC. Models besides Vicuna have been alignment-tuned and are already robust to prompts from JBC but vulnerable to other attacks. We find perplexity filtering is highly effective on GCG but is not effective on the natural language jailbreak prompts in PAIR and JBC. SmoothLLM is more effective across multiple attacks due to not relying on the perplexity heuristic. Still, it cannot defend against a significant proportion of prompts from PAIR, the strongest attack. Rephrasing is surprisingly effective, especially on JBC for Vicuna, outperforming the other defenses. We observe RPO is more effective than all baseline defenses on PAIR for all models, reducing ASR by 66% on Vicuna and 44% on GPT-4 and improving on the state-of-the-art defense SmoothLLM by 31% and 19%. This also shows that RPO suffixes transfer across models, as the suffix was optimized using Llama-2 but can transfer to Vicuna and even closed-source GPT models. Notably, RPO reduces GCG ASR to 0% for all models, fully defending against the attack. Using RPO with Llama-2 makes the model robust to all three attacks, the first time a model is fully robust on JailbreakBench. The only setting where RPO is not the strongest defense is JBC on Vicuna, where other defenses are more effective. This may be due to the lack of safety training on the older Vicuna model, making it less responsive to our defense.
Transfer to Unknown Attacks on HarmBench. HarmBench marks a difficult distribution shift from the optimization setup of RPO as it contains many attacks RPO was not optimized on and has a broader inclusion of behaviors, such as copyright infringement and contextual behaviors referencing user context. These categories are not covered in the instructions we optimize RPO on, forcing the defense to generalize. Despite this, in Tab. 2 we observe RPO transfers to all attacks in HarmBench for all models, generalizing to difficult attacks such as TAP and AutoDAN. Notably, RPO reduces ASR by an average of 18%, 6.6%, 8.7%, and 3.5% for Vicuna, Llama-2, GPT-3.5, and GPT-4, respectively. This suggests RPO is universally effective as a defense for harmful queries, irrespective of the attack. This is due to the defense enhancing existing refusal mechanisms in LLMs, which naturally transfers to other safety scenarios. However, we observe a much lower improvement on HarmBench than JailbreakBench, reflecting the challenges of generalizing to new attacks and behaviors.